Analyze the function f(x) = - 2 cot 3x. Include:
- Domain and range
- Period
- Two Vertical Asymptotes
1
Expert's answer
2020-12-20T13:41:58-0500

"domain\ :\ x \ne \dfrac{\pi}{6} + \dfrac{\pi n}{3} , \ n \in N\\
range \ :\ y \in R\\
period\ : \ \dfrac{\pi}{|b|} = \dfrac{\pi}{|3|} = \dfrac{\pi}{3}\\
 vertical\ asymptotes\ : x = \dfrac{\pi}{6} + \dfrac{\pi n}{3}, \ n \in N\\"